# Robustness of quantum algorithms: Worst-case fidelity bounds and implications for design

### v1

@@ D.2 Stochastic analysis from [34]
This bound indicates a favorable scaling of the error level with the circuit depth, i.e., $\delta\sim\frac{1}{\sqrt{N}}$ is sufficient for high fidelity, whereas the bound in Theorem II.1 results in $\delta\sim\frac{1}{N}$. However, it is important to emphasize that the bound (84) is only shown in [34] for the average fidelity and under independent and normally distributed noise. On the contrary, our results address arbitrary (bounded) noise and worst-case fidelity bounds, which are commonly assumed for threshold theorems [1, 35, 36, 37]. In particular, a scaling of $\delta\sim\frac{1}{N}$ can be expected for a general worst-case bound against coherent errors and without further structural assumptions since it cannot be excluded that the errors act in the same direction, which causes them to accumulate linearly.
@@
Appendix E Proof of Theorem III.1
@@
where $\bar{\mathcal{R}}_{N}=e^{N\mathcal{G}}$. The first term on the right-hand side can be bounded in complete analogy to the proof of Theorem II.1
@@
see [60, equation (5)]. Inserting this into (96), we infer
@@
where we use that the matrix $2$-norm is induced by the vector $2$-norm in the last inequality. Using straightforward algebraic manipulations analogous to the setup in Section II, one can prove that $\mathcal{A}=\bar{\mathcal{A}}\mathcal{R}_{N}$. Hence, the deviation of the noisy operation $\mathcal{A}$ from the ideal operation $\bar{\mathcal{A}}$ can be quantified by the deviation of $\mathcal{R}_{N}$ from identity. In particular, we have
@@
Similar to the proof of Theorem II.1, we have
@@
Appendix F Details on the QFT transpilation experiment
@@ F.1 QFT Circuits
Figure 7 shows the circuit diagram of untranspiled three-qubit textbook QFT circuit. We transpile the three-qubit textbook circuit into elementary base sets using the Berkeley Quantum Synthesis Toolkit (BQSKit) [68] with optimization level 2. The transpiled circuits for gate sets A-D results in the circuits shown in Figures 8-11.
@@
F.2 Details on numerically finding $\gamma$ via Appendix B.2
To determine the parameter $\gamma$ via the reformulation as an optimization problem (Appendix B.2), we employ the L-BFGS-B algorithm [75] as implemented in scipy. The optimization is initialized from 1000 randomly chosen starting points, and we select the largest $\gamma$ value obtained across all runs for further analysis. We verified that the algorithm converges reliably in all cases. Although we cannot guarantee that the global maximum is always found, the observed variation between the largest values obtained is marginal, providing strong evidence that the resulting $\gamma$ values are close to the global optimum.

### v2

@@ D.2 Stochastic analysis from [34]
This bound indicates a favorable scaling of the error level with the circuit depth, i.e., $\delta\sim\frac{1}{\sqrt{N}}$ is sufficient for high fidelity, whereas the bound in Theorem 2.1 results in $\delta\sim\frac{1}{N}$. However, it is important to emphasize that the bound (87) is only shown in [34] for the average fidelity and under independent and normally distributed noise. On the contrary, our results address arbitrary (bounded) noise and worst-case fidelity bounds, which are commonly assumed for threshold theorems [1, 38, 39, 40]. In particular, a scaling of $\delta\sim\frac{1}{N}$ can be expected for a general worst-case bound against coherent errors and without further structural assumptions since it cannot be excluded that the errors act in the same direction, which causes them to accumulate linearly.
@@
Appendix E Proof of Theorem 3.1
@@
where $\bar{\mathcal{R}}_{N}=e^{N\mathcal{G}}$. The first term on the right-hand side can be bounded in complete analogy to the proof of Theorem 2.1
@@
see [59, equation (5)]. Inserting this into (99), we infer
@@
where we use that the matrix $2$-norm is induced by the vector $2$-norm in the last inequality. Using straightforward algebraic manipulations analogous to the setup in Section 2, one can prove that $\mathcal{A}=\bar{\mathcal{A}}\mathcal{R}_{N}$. Hence, the deviation of the noisy operation $\mathcal{A}$ from the ideal operation $\bar{\mathcal{A}}$ can be quantified by the deviation of $\mathcal{R}_{N}$ from identity. In particular, we have
@@
Similar to the proof of Theorem 2.1, we have
@@
Appendix F Details on the numerical experiments
@@ F.1 QFT Circuits
Figure 6 shows the circuit diagram of untranspiled three-qubit textbook QFT circuit. We transpile the three-qubit textbook circuit into elementary base sets using the Berkeley Quantum Synthesis Toolkit (BQSKit) [67] with optimization level 2. The transpiled circuits for gate sets A-D results in the circuits shown in Figures 7-10.
@@
F.2 Modular adder circuit
The modular adder circuit used in the numerical experiment of this paper is based on [70]. We use the qiskit [77] implementation of this circuit and transpile this circuit into gate set B with the transpile method in qiskit.
F.3 Details on numerically finding $\gamma$ via Appendix B.2
To determine the parameter $\gamma$ via the reformulation as an optimization problem (Appendix B.2), we employ the L-BFGS-B algorithm [78] as implemented in scipy. The optimization is initialized from 100 randomly chosen starting points, and we select the largest $\gamma$ value obtained across all runs for further analysis. We verified that the algorithm converges reliably in all cases. Although we cannot guarantee that the global maximum is always found, the observed variation between the largest values obtained is marginal, providing strong evidence that the resulting $\gamma$ values are close to the global optimum.
Appendix G Effect of the partition size $\gamma$
The tightness of the bound on $\gamma$ is highly sensitive to the partition size and the specific bounding method employed. Generally, increasing the partition size (i.e., the maximum number of qubits in each sub-circuit) reduces the total number of partitions, thereby decreasing the conservatism introduced by the sequential bounding in (78). However, this comes at the cost of computational complexity, which scales exponentially with the number of qubits.
Figure 11: The bound on $\gamma_{\mathrm{norm}}$ for a 50 qubit modular adder where the circuit is partitioned into smaller sub-circuits with different partition sizes.
When calculating the norm-based bound $\gamma_{\mathrm{norm}}$ (Appendix B.4), an additional competing effect arises from the Frobenius norm used in the inequalities of (75). This causes the bound to scale with the number of qubits involved in the circuit. By maintaining small partition sizes, the conservatism introduced by these norm inequalities remains manageable relative to that of the partition-based summation. This trade-off is illustrated in Figure 11, which depicts the $\gamma_{\mathrm{norm}}$ bound for the 50-qubit modular adder across various partition sizes. The data demonstrates that for this specific method, smaller partition sizes can lead to less conservative estimates.